# CoRe: Context-Robust Remasking for Diffusion Language Models

### v2

@@ Appendix E Qualitative Examples
Fig. 5 shows a math problem comparison between CoRe and ReMDM on the GSM8K dataset. Given the prompt, our calculation expects the correct result to be 51. In the CoRe example (left), the model performs the intermediate reasoning correctly and computes the value “51”. However, it still answers “151” due to an early commitment toward an extra token “1”. CoRe detects the discrepancy on step 40, and remasks the leading digit “1” which is replaced by a space in the same step, yielding an answer that is consistent with the question. On the other hand, ReMDM-conf (right) fails to detect this same artifact, leaving the incorrect output intact across subsequent steps.
@@
Figure 5: CoRe fixes error in output format. CoRe corrects the wrong answer “151” to the correct “51” by replacing the token “1” with a space. In contrast, ReMDM-conf focuses on the token <|endoftext|> which is unrelated to the error, and fails to correct the underlying mistake.
Fig. 6 presents a MATH example illustrating how CoRe can gradually correct an initial mistake in its reasoning trajectory. In step 48, we observe the base LLaDA model has generated the sequence “x = =.”. CoRe detects the syntax error in the mathematical expression, and revises the equal sign to “1”. Then, although the base model reasons correctly, it produces the wrong answer of “-6”, while our visual inspection suggests that the answer should be “-2”. Again, CoRe detects the error in step 88, which is fixed in step 89. On the other hand, ReMDM-conf repeatedly fails to focus on the context-brittle tokens, and ends with a flawed answer.
@@
Figure 6: CoRe revises reasoning trajectory. CoRe first removes an incorrect syntax in the reasoning which is inconsistent with the later derivation, and in a subsequent step revises the actual answer “-2” which the sampler mistakenly wrote as “-6”. These staged revisions allow the solution to converge to the correct final answer. In contrast, ReMDM-conf revises nearby but non-erroneous tokens, leaving the incorrect intermediate computation unchanged and resulting in an incorrect final answer. (Ellipses (...) denote code segments that remain unchanged across steps and are omitted for brevity.)
@@
Fig. 8 compares CoRe with ReMDM-conf on another programming example where the cause is a corrupted function signature. Instead of generating “opposite_Signs(num1, num2)”, the base model generated “opposite_Signs(num1, :)2)”. The standard LLaDA sampler would not be able to address this issue. In step 32, CoRe detects the context-brittle “:)” token and revises it to “num”, yielding a coherent program. In contrast, ReMDM-conf repeatedly masks tokens that are not responsible for the failure, hence the tokens gets sampled to the original values. Together, these examples demonstrate that CoRe consistently directs revisions toward tokens that constrain the evolving structure of the solution, whereas ReMDM-conf often focuses on irrelevant tokens.
@@
Figure 8: CoRe recovers corrupted function signatures. CoRe identifies context-brittle tokens whose likelihood collapses under perturbation of the surrounding context. As a result it correctly identifies the erroneous symbols, which are revised to form the variable name “num2”, yielding a structurally coherent program that can pass the given assertations. In contrast, ReMDM-conf keeps trying to focus on tokens that are not erroneous, resulting in a code that is structurally wrong according to Python syntax.

### v3

@@ Appendix E Qualitative Examples
Fig. 5 shows a math problem comparison between CoRe and ReMDM-conf on the GSM8K dataset. Given the prompt, the expected result is 51. In the CoRe example (left), the model performs the intermediate reasoning as intended and computes the value 51. However, it still answers 151 due to an early commitment toward an extra token 1. CoRe detects the discrepancy on step 40 and remasks the leading digit 1, which is replaced by a space in the same step, yielding an answer that is consistent with the prompt. On the other hand, ReMDM-conf (right) fails to detect this same artifact, leaving the inconsistent output intact across subsequent steps.
@@
Figure 5: CoRe resolves inconsistencies in output format. CoRe revises the answer 151 to 51 by replacing the token 1 with a space. In contrast, ReMDM-conf focuses on the <|endoftext|> token, which is unrelated to the error, and fails to resolve the underlying issue.
Fig. 6 presents a MATH example illustrating how CoRe can gradually correct an initial mistake in its reasoning trajectory. In step 48, we observe that the base LLaDA model has generated the sequence x = =.. CoRe detects the syntax error in the mathematical expression and revises the equal sign to 1. Then, although the base model reasons correctly, it produces the wrong answer of -6, while our visual inspection suggests that the answer should be -2. Again, CoRe detects the error in step 88, which is fixed in step 89. On the other hand, ReMDM-conf repeatedly fails to focus on these context-brittle tokens, ultimately yielding an incorrect answer.
@@
Figure 6: CoRe revises reasoning trajectory. CoRe first removes an incorrect syntax in the reasoning which is inconsistent with the later derivation, and in a subsequent step revises the actual answer -2 which the sampler mistakenly wrote as -6. These staged revisions allow the solution to converge to the correct final answer. In contrast, ReMDM-conf revises nearby but non-erroneous tokens, leaving the incorrect intermediate computation unchanged and resulting in an incorrect final answer. (Ellipses (...) denote code segments that remain unchanged across steps and are omitted for brevity.)
@@
Fig. 8 compares CoRe with ReMDM-conf on another programming example where the cause is a corrupted function signature. Instead of generating opposite_Signs(num1, num2), the base model generated opposite_Signs(num1, :)2). The standard LLaDA sampler would not be able to address this issue. In step 32, CoRe detects the context-brittle :) token and revises it to num, yielding a coherent program. In contrast, ReMDM-conf repeatedly masks tokens that are not responsible for the failure, hence the tokens gets sampled to the original values. Together, these examples demonstrate that CoRe consistently directs revisions toward tokens that constrain the evolving structure of the solution, whereas ReMDM-conf often focuses on irrelevant tokens.
@@
Figure 8: CoRe recovers corrupted function signatures. CoRe identifies context-brittle tokens whose likelihood collapses under perturbation of the surrounding context. As a result it correctly identifies the erroneous symbols, which are revised to form the variable name num2, yielding a structurally coherent program that can pass the given assertations. In contrast, ReMDM-conf keeps trying to focus on tokens that are not erroneous, resulting in a code that is syntactically incorrect.